OpenStudy (anonymous):

Simplify the expression: -20a ^{-2}b ^{-7} / 5a ^{-5}b ^{3} Write your answer without negative exponents.

OpenStudy (anonymous):

\[-20a ^{-2}b ^{-7} / 5a ^{-5}b ^{3}\] There is the equation how it is supposed to look.

OpenStudy (skullpatrol):

$$\frac{-20a ^{-2}b ^{-7}}{ 5a ^{-5}b ^{3}}$$

OpenStudy (anonymous):

Yeah, thats the equation. I'm just a little lost.

OpenStudy (skullpatrol):

The first thing you should understand is this is not an "equation," it is an "expression" do you know why?

OpenStudy (anonymous):

Because it doesn't have an equals sign to equal a value.

OpenStudy (skullpatrol):

Correct :) An "equation" must have an "equals sign"

OpenStudy (anonymous):

So when I'm dividing two numbers like this, don't I divide all whole number and subtract like variables?

OpenStudy (skullpatrol):

There are certain "rules of exponents" that you should understand first. You are perfectly correct in working with "like variables" by this I mean the bases must be the same.

OpenStudy (anonymous):

So I can subtract the exponents on say \[a^{-2}\] and \[a ^{-5}\]?

OpenStudy (skullpatrol):

Yes, -3-(-5) is the exponent on a.

OpenStudy (anonymous):

But the first A has a negative two on it?

OpenStudy (anonymous):

And then that gives us -7, but the question says to write the answer without negative exponents. This is where I start to get lost...

OpenStudy (skullpatrol):

Ooops I meant -2-(-5) is the exponent on a.

OpenStudy (anonymous):

Okay so that gives us A^-7, and like I said... the original question asks to write it without negative exponents. I kinda get lost whenever I come to this point.

OpenStudy (skullpatrol):

What is -2-(-5)=?

OpenStudy (anonymous):

OH! its a negative negative... so the answer to that would be 3?

OpenStudy (skullpatrol):

Yes :)

OpenStudy (anonymous):

Okay, and just one final question. On the -20 / 5... its just that, right? -20 / 5?

OpenStudy (anonymous):

Actually one more question now that I look at that. The b^-7 and b^3 would be -7 + 3, correct?

OpenStudy (skullpatrol):

How many times does 5 go into 20?

OpenStudy (anonymous):

4

OpenStudy (skullpatrol):

So -20 / 5 =?

OpenStudy (anonymous):

-4

OpenStudy (skullpatrol):

Correct :)

OpenStudy (anonymous):

Okay, now one more question (sorry). The -7 + 3 on the B's would be -4. What am I not seeing there?

OpenStudy (skullpatrol):

The exponent is 10 on b and it belongs in the denominator.

OpenStudy (anonymous):

How?

OpenStudy (skullpatrol):

3-(-7)

OpenStudy (anonymous):

Ah okay, so it would be \[-4a ^{3}b ^{10}\]

OpenStudy (skullpatrol):